40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1419     Entry time: Tue Mar 24 03:05:25 2009
Author: Yoichi 
Type: Update 
Category: Locking 
Subject: Locking tonight 
MC1 issue:
The MC1 seems to be drifting still. I found it was off from the SUS drift-mon reference values and restored the alignment using the SUS drift-mon before I went home for dinner.
But when I came back being happy with the Japanese victory over S-Korea at the WBC final, the MC was unhappy again.
I restored the alignment of the MC1 using the SUS drift-mon once again and centered the WFS QPDs.
I will leave the MC unlocked again tonight to see the drift. You are welcome to lock the MC in the morning as I will have corrected enough data by the time people come in.

Computer overloads:
I removed some filters from suspensions to off load susvme computers.
Nonetheless, both susvme1 and susvme2 are still over loaded during the dither alignment. The alignment results are in general ok. So this is not a too serious problem.
But still it would be nice to resolve.

3.8kHz hunting:
I made several measurements of the AO path loop gains (using the SR785) and the transfer functions from the CARM excitation (actuation to the ETMs) to the PO_DC signal as the arm powers are increased.
There is a similar structure as in the AO loop found also in the CARM->PO_DC transfer functions. This implies that the problem is likely to be in the PO_DC sensor not in the MC->VCO actuator. But the MC and the VCO could still be the
cause of the problem because they were in the control loop when the CARM->PO_DC TF were measured.
The peak frequency does not seem to depend on the arm power, but the conclusion is not definite because I was only able to measure the TFs from arm power 5 to 10 (not much difference).
I will make plots and post them later.

To Do for tomorrow:
Tonight the CARM error signal was noisier than the reference spectra (broad band white noise appeared). I should check the beam centering of the SPOB PD.
Also someone should center the oplevs of the mirrors as some of them are off.
Continue to measure the TFs at various power levels.
Try to put another (Thorlabs?) PD at the POB port to get PO_DC from it.
ELOG V3.1.3-